OpenStudy (anonymous):

can anyone help me how to find r1 and r2 of this qustion We are given an input voltage V, and we need to provide an open-circuit output voltage of V. An additional requirement is that the Thevenin resistance as seen from the output terminals is between and . Assume first that the resistors have their nominal resistance. Come up with resistors and such that the division ratio is within 10% of the requirement. Of course, the resistances you chose are just nominal. Given that they are only guaranteed to have resistances within 10% of the nominal value, what is the larg

OpenStudy (anonymous):

username "00585711" asked this and ali110 answered it for certain variables (which sure helped me out)... but i guess i can try to give u a start. first the problem gives u a Vin and Vout. Vout/Vin =ratio step 2 u must bound your ratio by +/- 10% in order to find your R1. For example... Vout=24 and Vin = 80 then your ratio is 0.3. 0.3*.1= 0.03 so 0.3+0.03=0.33(upper) && 0.3-0.03=0.27(low) plug in your upper/lower bound to squeeze your R1. \[\frac{ Rthev1 }{ upper } <R1<\frac{ Rthev2 }{ low }\] where Rthev1=lower and Rthev2=higher should be given. you should now have a range of values that can be rounded off nicely and correspond to the set E12 given in the question. set one of the E12 values as R1 then use your \[\frac{ Vout }{ Vin }=\frac{ R2 }{ R2+R1 }\] and solve for R2 and again round it off to the nearest value given by the E12 set. hopefully this helps get u started... if u need more help, ali110 already did a very good job of explaining this problem.